# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ Abstract
Topological data analysis (TDA) is a rapidly growing area of data science which uses the geometry and topology of data sets to produce qualitative multi-scale shape descriptors for subsequent statistical and machine learning tasks. The most common descriptor in TDA is persistent homology, which tracks the topological changes in growing families of subsets of the data set itself, called filtrations, and encodes them in an algebraic object, called a persistence module. The algorithmic and theoretical properties of persistence modules are now well understood in the single-parameter case, that is, when there is only one filtration (e.g., feature scale) to study. In contrast, much less is known in the multi-parameter case, where several filtrations (e.g., scale and density) are used simultaneously. Indeed, the resulting multi-parameter persistence modules are much more complicated and intricate, which dramatically impedes the study of their theoretical properties. However, they usually encode information that is invisible to their single-parameter counterparts, and are thus much more useful descriptors for applications in data science. As a consequence, a lot of attention has been devoted to the construction of tractable and stable proxies for multi-parameter persistence modules. However, most of the proposed approaches in the literature are still prohibitively expensive to compute on large-scale data, many are limited to at most two filtrations, and the most tractable sacrifice much of the richness of the multi-parameter setting.
In this article, we introduce a new parameterized family of topological invariants, taking the form of candidate decompositions, for multi-parameter persistence modules. We prove that our candidate decompositions are controllable approximations: when restricting to modules that can be decomposed into interval summands, we establish theoretical results about the approximation error between our candidate decompositions and the true underlying module in terms of the standard interleaving and bottleneck distances. Moreover, even when the underlying module does not admit such a decomposition, our candidate decompositions are nonetheless stable invariants; small perturbations in the underlying module lead to small perturbations in the candidate decomposition. Then, we introduce MMA (Multipersistence Module Approximation): an algorithm for computing stable instances of such invariants, which is based on fibered barcodes and exact matchings, two constructions that stem from the theory of single-parameter persistence. By design, MMA can handle an arbitrary number of filtrations, and has bounded complexity and running time. Finally, we present empirical evidence validating the generalization capabilities and running time speed-ups of MMA on several data sets.
@@ 1 Introduction
Topological Data Analysis (TDA) [EH10, Oud15] is a new and rapidly developing area of data science that has seen a lot of interest due to its success in various applications, ranging from bioinformatics [RB19] to material science [BHO18]. The main computational tool of TDA is persistent homology (PH). Whereas homology is a qualitative descriptor of the shape of a topological space $S$ [^1] In this article, we work with simplicial homology, hence it is assumed throughout the article that $S$ is a simplicial complex., the core idea of PH is to capture how the homology groups change when computed on a filtration of $S$. A filtration is a family $\{S_{x}\subseteq S\}_{x\in\mathcal{A}}$ of subspaces of $S$ indexed over a partially ordered set (poset) $\mathcal{A}$, that is nested w.r.t. inclusion, i.e., it satisfies $S_{x}\subseteq S_{y}$ for any $x\leq y$. Then, the functoriality of homology and these inclusions induce morphisms between the corresponding homology groups $H_{*}(S_{x})\rightarrow H_{*}(S_{y})$ for each pair $x\leq y$, which allows to detect the differences in homology when going from index $x$ to index $y$. One of the most common ways to produce such filtrations is to study the sublevel sets of a continuous filter function $f:S\rightarrow\mathbb{R}^{n}$, defined with $S_{x}=f^{-1}(\{x^{\prime}\in\mathbb{R}^{n}:x^{\prime}\leq_{n}x\})$; the partial order on the poset $\mathbb{R}^{n}$ (denoted by $\leq_{n}$) is defined, for $x,y\in\mathbb{R}^{n}$, as $x\leq_{n}y$ if and only if $x_{i}\leq y_{i}$ for every dimension $i$.
@@
When $\mathcal{A}$ is totally ordered, e.g., when $\mathcal{A}\subseteq\mathbb{R}$, then applying the homology functor $H_{*}(-;k)$ for a field $k$ to a (single-parameter) filtration results in a sequence of vector spaces connected by linear maps, called a single-parameter persistence module. This situation has been studied extensively in the TDA literature [Car09, CdSGO16, EH10, Oud15]. Notably, one can show that such persistence modules can always be decomposed into a direct sum of simple interval summands: ${\mathbb{M}}\simeq\bigoplus_{i\in\mathcal{I}}I(b_{i},d_{i})$, where each interval summand $I(b_{i},d_{i})$ intuitively represents the lifetime of a topological structure, i.e., $b_{i}$ is the appearance time (birth) and $d_{i}$ is the disappearance time (death) of a topological structure, that is detected by homology as the index increases. Moreover, single-parameter persistence modules can be efficiently represented in a compact descriptor called the persistence barcode, and several representation methods, such as Euclidean embeddings and kernels for machine learning classifiers, have been proposed for such barcodes in the literature [Bub15, ${}^{+}$17, RHBK15, CCO17, ${}^{+}$20]. As a consequence, most applications of TDA use single-parameter persistence modules, and often use the sublevel sets of, e.g., the data set scale, as the corresponding single-parameter filtration.
@@
However, many data sets come with not just one, but multiple, possibly intertwined, salient filtrations. For example, image data typically has both a spatial filtration and an intensity filtration, and arbitrary point cloud data can be filtered both by feature scale and density. Unfortunately, in general, the resulting multi-parameter persistence modules, obtained by applying the homology functor to a filtration indexed over $\mathbb{R}^{n}$ [BL22], are much less tractable; in contrast to the single-parameter case, there is no decomposition theorem that can break down any module into a direct sum of simple indicator summands (e.g., interval modules). Instead, there is now a rich literature on general decompositions into arbitrarily complicated summands [DX22] and their associated minimal presentations [KR21, LW19], and on the theoretical study of a few restricted cases (such as $n=2$ filtrations or multi-parameter persistence modules in homology dimension 00) where simple decompositions can be obtained [${}^{+}$18, AENY19, BLOry, CO19, DX21]. It has thus become crucial to define general topological invariants for multi-parameter persistence modules that are meaningful, visually interpretable, and easily computable.
@@
In this article, we introduce new invariants of multi-parameter persistence modules (obtained from the persistent homology of multi-parameter filtrations of simplicial complexes), along with a new algorithm we call MMA for their practical computations. More precisely, our contributions are five-fold:
@@
    We introduce a new family of topological invariants for multi-parameter persistence modules (Definition 3.3), taking the form of candidate decompositions $\tilde{{\mathbb{M}}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}\tilde{I}_{i}$. These candidate decompositions are parameterized by a precision parameter $\delta>0$, and each $\tilde{I}_{i}$ in these candidate decompositions is an interval summand in $\mathbb{R}^{n}$.
@@
    When computed over interval decomposable modules, we prove that under mild assumptions the interleaving and bottleneck distances between our candidate decompositions $\tilde{\mathbb{M}}_{\delta}$ and the module ${\mathbb{M}}$ they approximate are upper bounded (Proposition 3.4):
    ${d_{\rm I}}({\mathbb{M}},\tilde{\mathbb{M}}_{\delta})\leq{d_{\rm B}}({\mathbb{M}},\tilde{\mathbb{M}}_{\delta})\leq\delta.$
@@
    We introduce our method MMA (Multi-parameter persistence Module Approximation, Algorithm 1) for computing instances of such candidate decompositions, with running time
@@
  4. 4.
    We study the approximation and stability properties of MMA. More precisely, we show that when $\delta$ becomes small enough, the candidate decompositions $\tilde{\mathbb{M}}_{\delta}^{\texttt{MMA}{}}$ computed by MMA recover the underlying persistence module ${\mathbb{M}}$ exactly, provided that it is interval decomposable (Corollary 4.10):
    ${d_{\rm I}}({\mathbb{M}},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})={d_{\rm B}}({\mathbb{M}},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})=0,$
    when $\delta$ is less than a constant that depends only on the multi-parameter filtration values. In the general case (when ${\mathbb{M}}$ is not necessarily interval decomposable), we show that the candidate decompositions produced by MMA preserve the (single-parameter) persistence barcodes associated to diagonal slices of the multi-parameter filtration (Proposition 4.12):
    ${d_{\rm B}}\left({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}),{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}\vphantom{\big{|}}\right|_{l}})\right)\leq 2\delta,$
    and that they are also stable w.r.t. the input data (Proposition 4.13):
@@
    We discuss practical choices of the input parameters of MMA in Section 5, and we perform numerical experiments that showcase the performance of MMA and exhibit the trade-off between computation time and approximation error in Section 6.
@@
An important input parameter of our method MMA is the so-called matching function $\sigma$, which has a large influence on the candidate decompositions produced by MMA. While the output of MMA is only an approximation of the underlying module ${\mathbb{M}}$, we hypothesize that, in the general case, the family $\mathcal{F}$ of candidate decompositions obtained from MMA by varying the matching function parameter across an appropriate family of matching functions $\Sigma$, as =$\mathcal{F}=\{\texttt{MMA}{}({\mathbb{M}},L_{\delta},\sigma)\}_{\sigma\in\Sigma,\delta>0}$ , is a complete invariant of the underlying module ${\mathbb{M}}$. We leave this conjecture open for future work.
@@
The most common topological invariants of multi-parameter persistence modules are the so-called Hilbert function ${\rm HF}$ and rank invariant ${\rm RI}$. The Hilbert function simply gathers the dimensions of all vector spaces in the module ${\rm HF}=x\mapsto{\rm dim}(H_{*}(S_{x}))$, i.e., it counts the number of topological structures that are present in $S_{x}$ at index $x$, while the rank invariant gathers the ranks of all morphisms in the module ${\rm RI}=(x,y)\mapsto{\rm rank}(H_{*}(S_{x})\to H_{*}(S_{y}))$ (if $x\leq_{n}y$), i.e., it counts the topological structures that are preserved when going from index $x$ to index $y$. The rank invariant is equivalent to computing the persistence barcodes associated to all possible lines in $\mathbb{R}^{n}$; this slicing process is usually referred to as the fibered barcode. The Hilbert function, the rank invariant and the fibered barcode can all be computed with the Rivet software [LW15] in $O(N^{3}\kappa+(N+{\rm log}\kappa)\kappa^{2})$ operations, where $N$ is the number of simplices, and $\kappa$ is the number of indices on which ${\rm HF}$ and ${\rm RI}$ are evaluated. Note that Hilbert functions and rank invariants can also be decomposed themselves in order to be more visually interpretable—the corresponding invariants are called signed barcodes [BOO21].
Unfortunately, Rivet is currently limited to at most $n=2$ filtrations. Moreover, invariants based on dimensions and ranks are known to miss a lot of the structural properties of modules: it is quite easy to find examples of distinct modules that share the same Hilbert function and rank invariant [Vip20a, Figure 3]. Thus, generalizations of the Hilbert function, the rank invariant, the signed barcodes, and their corresponding algorithms, have been introduced recently [DKM22, KM18]. To our knowledge, there is however no corresponding available implementations of these tools.
Finally, another recent line of work also provides candidate decompositions into simple summands for modules that are computed from point cloud data. These decompositions are obtained from applying the elder rule and called the elder-rule staircodes [CKMW20], with running time $O(m^{2}\log(m))$ where $m$ is the number of points of the point cloud. This approach has been shown to be to be theoretically close to the true decomposition, but only works in homology dimension 0 and for $n=2$ parameters.
Hence, computing candidate decompositions for multi-parameter persistence modules indexed over $\mathbb{R}^{n}$ for arbitrary $n\in\mathbb{N}^{*}$ with controlled complexity, running time, and approximation error is still an open and important question, which we tackle in this article using our approximation scheme MMA.
@@
Section 2 provides a concise review of multi-parameter persistence modules. In Section 3, we present our new invariants for multi-parameter persistence modules and their approximation properties. Then, we introduce our corresponding algorithm MMA and its stability properties in Section 4. We also discuss the choice of its input parameters in Section 5. Finally, we illustrate the performances of MMA in Section 6.

### v4

# Multi-parameter Module Approximation: an efficient and interpretable invariant for multi-parameter persistence modules with guarantees

@@ Abstract
Topological data analysis (TDA) is a rapidly growing area of data science, whose most common descriptor is persistent homology, which tracks the topological changes in growing families of subsets of the data set itself, called filtrations, and encodes them in an algebraic object, called a persistence module. The algorithmic and theoretical properties of persistence modules are now well understood in the single-parameter case, that is, when there is only one filtration (e.g., feature scale) to study. In contrast, much less is known in the multi-parameter case, where several filtrations (e.g., scale and density) are used simultaneously. Since multi-parameter persistence modules usually encode information that is invisible to their single-parameter counterparts, it is critical to build tractable proxies for them, ideally with some theoretical robustness guarantees.
In this article, we introduce a new parameterized family of topological descriptors, taking the form of candidate decompositions, for multi-parameter persistence modules, and we a identify a subfamily of these descriptors, that we call approximate decompositions, that are controllable approximations, in the sense that they preserve diagonal barcodes. Then, we introduce MMA (Multipersistence Module Approximation): an algorithm based on matching functions for computing instances of candidate decompositions with some precision parameter $\delta>0$. By design, MMA can handle an arbitrary number of filtrations, and has bounded complexity and running time. Moreover, we prove the robustess of MMA: when computed with so-called compatible matching functions, we show that MMA produces approximate decompositions (and we prove that such matching functions exist for $n=2$ filtrations). Next, we restrict the focus on modules that can be decomposed into interval summands. In that case, compatible matching functions always exist, and we show that, for small enough $\delta$, the approximate decompositions obtained with such compatible matching functions by MMA have an approximation error (in terms of the standard interleaving and bottleneck distances) that is bounded by $\delta$, and that reaches zero for an even smaller, positive precision $\delta_{\rm exact}$. Finally, we present empirical evidence validating that MMA has state-of-the-art performance and running time on several data sets.
Keywords: Multiparameter Persistent Homology, Persistence Modules, Interval Modules, Approximation Methods, Convergence Analysis
@@ 1 Introduction
Topological Data Analysis (TDA) (35, 54) is a new and rapidly developing area of data science that has seen a lot of interest due to its success in various applications, ranging from bioinformatics (55) to material science (16). The main computational tool of TDA is persistent homology (PH). Whereas homology is a qualitative descriptor of the shape of a topological space $S$, the core idea of PH is to capture how the homology groups change when computed on a filtration of $S$. A filtration is a family $\{S_{x}\subseteq S\}_{x\in A}$ of subspaces of $S$ indexed over a partially ordered set (poset) $A$, that is nested w.r.t. inclusion, i.e., it satisfies $S_{x}\subseteq S_{y}$ for any $x\leq y$. Then, the functoriality of homology and these inclusions induce morphisms between the corresponding homology groups $H_{*}(S_{x})\rightarrow H_{*}(S_{y})$ for each pair $x\leq y$, which allows to detect the differences in homology when going from index $x$ to index $y$. One of the most common ways to produce such filtrations is to study the sublevel sets of a continuous filter function $f:S\rightarrow\mathbb{R}^{n}$, defined with $S_{x}=f^{-1}(\{x^{\prime}\in\mathbb{R}^{n}:x^{\prime}\leq_{n}x\})$; the partial order on the poset $\mathbb{R}^{n}$ (denoted by $\leq_{n}$) is defined, for $x,y\in\mathbb{R}^{n}$, as $x\leq_{n}y$ if and only if $x_{i}\leq y_{i}$ for every dimension $i$.
@@
When $\mathcal{A}$ is totally ordered, e.g., when $\mathcal{A}\subseteq\mathbb{R}$, then applying the homology functor $H_{*}(-;k)$ for a field $k$ to a (single-parameter) filtration results in a sequence of vector spaces connected by linear maps, called a single-parameter persistence module. This situation has been studied extensively in the TDA literature (18, 24, 35, 54). Notably, one can show that such persistence modules can always be decomposed into a direct sum of simple interval summands: ${\mathbb{M}}\simeq\bigoplus_{i\in\mathcal{I}}{k^{I(b_{i},d_{i})}}$, where each interval summand $k^{I(b_{i},d_{i})}$ intuitively represents the lifetime $I(b_{i},d_{i})$ of a topological feature, i.e., $b_{i}$ is the appearance time (birth) and $d_{i}$ is the disappearance time (death) of a topological feature, that is detected by homology as the index increases. Moreover, single-parameter persistence modules can be efficiently represented in a compact descriptor called the persistence barcode, and several representation methods, such as Euclidean embeddings and kernels for machine learning classifiers, have been proposed for such barcodes in the literature (15, 1, 56, 22, 21). As a consequence, most applications of TDA use single-parameter persistence modules, and often use the sublevel sets of, e.g., the data set scale, as the corresponding single-parameter filtration.
@@
However, many data sets come with not just one, but multiple, possibly intertwined, salient filtrations. For example, image data typically has both a spatial filtration and an intensity filtration, and arbitrary point cloud data can be filtered both by feature scale and density. Unfortunately, in general, the resulting multi-parameter persistence modules, obtained by applying the homology functor to a filtration indexed over $\mathbb{R}^{n}$ (13), are much less tractable; in contrast to the single-parameter case, there is no decomposition theorem that can break down any module into a direct sum of simple indicator summands (e.g., interval modules). Instead, there is now a rich literature on general decompositions into arbitrarily complicated summands (34, 32) and their associated minimal presentations (36, 38, 48), on the theoretical study of a few restricted cases (such as some specific $n=2$-parameter filtrations or exact p.f.d. (Definition 2) $2$-parameter persistence modules) where simple decompositions can be obtained (27, 10, 4, 11, 44), and on simpler representations of multi-parameter persistence modules, such as the Euler characteristic (37), the fibered barcode (29, 57), the signed barcode (14, 51), and the (generalized) rank invariant and persistence diagram (41, 59). It has thus become crucial to define general topological descriptors for multi-parameter persistence modules that are meaningful, visually interpretable, and easily computable.
@@
In this article, we introduce new descriptors for multi-parameter persistence modules ( computable from multi-parameter filtrations of simplicial complexes [^2] Or even solely from module presentations, see Remark 13.) along with a new algorithm that we call MMA (Multipersistence Module Approximation) for their practical computations.
[…]
@@
    We fix a grid of diagonal lines $L$ spaced by $\delta>0$, and compute the barcodes associated to ${({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}})}_{l\in L}$.
@@
    Using some matching function $\sigma$, we match bars of consecutive barcodes together.
@@
[…]
Our contributions are five-fold:
  1. 1.
    We introduce a new family of topological descriptors for finitely presented multi-parameter persistence modules (Definition 19), taking the form of candidate decompositions $\tilde{\mathbb{M}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}{k^{\tilde{I}_{i}}}$. These candidate decompositions are parameterized by a precision parameter $\delta>0$, and each $k^{\tilde{I}_{i}}$} in these candidate decompositions is an interval summand in $\mathbb{R}^{n}$.
  2. 2.
    Then, we introduce our method MMA (Multi-parameter persistence Module Approximation, Algorithm 1) for computing instances of such candidate decompositions. Our method is crucially based on so-called matching functions, and, using any matching function whose complexity is linear w.r.t. $N$, has running time
@@
  3. 3.
    We show that MMA is a good approximation when computed with so-called compatible matching functions, i.e., that the candidate decompositions produced by MMA are, in this case, approximate decompositions: they preserve the (single-parameter) persistence barcodes associated to diagonal slices of the multi-parameter filtration (Proposition 3):
    $\textnormal{for all diagonal line }l\subseteq\mathbb{R}^{n},\quad{d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)\leq 2\delta.$
    We also show that, upon carefully choosing matching functions, the approximate decompositions produced by MMA are also stable w.r.t. the input data (Proposition 4):
@@
  4. 4.
    Then, we restrict the focus to interval decomposable modules. In that case, compatible matching functions always exist, and, under generic assumptions and small enough $\delta$, we prove that the interleaving and bottleneck distances between the approximate decompositions produced by MMA and the underlying persistence module are upper bounded (Proposition 5):
    ${d_{\rm I}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})\leq{d_{\rm B}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})\leq\delta.$
    Moreover, we also show that, when $\delta\leq\delta_{\rm exact}$, where $\delta_{\rm exact}$ is a constant that depends only on the multi-parameter filtration values, the approximate decompositions produced by MMA recover the underlying persistence module exactly (Proposition 12):
    ${d_{\rm I}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})={d_{\rm B}}(\mathbb{M},\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta})=0.$
@@
    We perform numerical experiments that showcase the performance of MMA and exhibit the trade-off between computation time and approximation error in Section 7.
@@
Computing approximate decompositions of multi-parameter persistence modules with simple summands or interval summands has been studied in a few other works. For instance, in (33), the authors provide an algorithm that computes optimal approximations with rectangle summands of interval decomposable, $2$-parameter persistence modules in order to lower bound their interleaving distance. In (5), the authors provide a method that associates to every $2$-parameter persistence module (and that was recently generalized to any $n$-parameter persistence module (6)) a pair of interval decomposable persistence modules obtained by computing the Möbius inversion of the so-called compressed multiplicity, in a spirit similar to the computation of generalized persistence diagrams and signed barcodes. Focusing on homology dimension 0 and $n=2$ with the Vietoris-Rips bifiltration on augmented metric spaces, the authors in (17) have proposed the elder-rule staircode, an interval decomposable persistence module who is known to recover the rank invariant and even the true interval decomposition of the module (if it exists) under some assumptions. Focusing rather on stability (as general decompositions of modules are known to be highly instable), in (7) the author provides new decompositions for persistence modules based on pruning (although with arbitrarily complicated summands) that enjoy better robustness guarantees.
More closely related to our approach, both (20) and (39, 40) provide methods to compute descriptors for $2$-parameter persistence modules using matching functions between persistence barcodes computed from a sorted family of $1$-dimensional slices, namely the multi-parameter persistence image, and the graphcode, respectively. While the multi-parameter persistence image encodes a decomposition constructed by matching consecutive barcodes with the vineyards algorithm (28) and computing the summand boundaries with the endpoints of matched bars (which does not guarantee that the resulting summand is an interval, or even that the resulting module is close to the input), the graphcode is an abstract graph whose nodes represent the bars of consecutive barcodes, and whose edges are built based on matching functions representing canonical inclusions of representative cycles, after fixing some cycle bases for every slice. Note that these matchings are not one-to-one: a representative cycle in one barcode might be included in several others in the next barcode.
Our method differs from the previous ones in three key aspects: first, note that, except for (6, 7), all other approaches are designed for $n=2$ filtrations, while our method MMA can handle an arbitrary number $n$ of filtrations (albeit without theoretical guarantees if $n>2$) and works in any homology dimension. Second, we designed MMA so that it has parameterized complexity: while the other approaches might be costly to compute as they rely on, e.g., Möbius inversions on large non-grid posets, the running time for MMA is controlled by the user through the choice of $\delta$ and of the matching function. Finally, and most importantly, our intention with the interval decomposable module produced by MMA was to provide a descriptor that is:
  1. (i)
    interpretable in the same way than persistence barcodes: its summands should correspond to lifetimes (in $\mathbb{R}^{n}$) of some homologous representative cycles, and
  2. (ii)
    as stable as possible: as it is impossible to provide interval decompositions that are always consistent with the rank invariant of the module (see (46, Section 10.2.3)), and as we still want to encode more information than the pointwise dimension of the module (i.e., the Hilbert function, which already requires a cubic complexity to compute in the two parameter case (26)), we seek for decompositions that preserve the persistence barcodes of diagonal slices of the module.
We prove in this article that our method MMA produces interval decompositions that satisfy both Items (i) and (ii). On the other hand, other approaches are either less interpretable (33, 5, 6, 7, 39, 40), thus not satisfying (i), or not stable enough (20, 17), thus not satisfying (ii).
More precisely, concerning (ii), our method can be seen as a generalization of the decompositions provided in (17), as we are not restricted to homology dimension 0, and as a continuation of the decompositions provided in (20), as $(a)$ we generalize it to compatible matching functions instead of relying solely on the vineyards algorithm, $(b)$ we guarantee that the produced decompositions only contain interval summands instead of producing summands supported on arbitrarily complicated shapes, and $(c)$ we prove that the produced decomposition preserve the diagonal barcodes, and recover the true decompositions exactly when the input module is itself interval decomposable.
Finally, one might argue that the interpretability property that we ask in (i) also induces some degree of arbitrariness, as there are, in general, several ways to choose representative cycles and their lifetimes (precisely because there are no good barcodes in multi-paramerer persistent homology, again see (46, Section 10.2.3)). This is perfectly true, and illustrates the trade-off between interpretability and computability that is often encountered in this field; note for instance that graphcodes (39, 40) also rely on arbitrary choices of cycles bases. However, notice that $(a)$ using such interval decompositions has nonetheless already proved to be useful in computational topology (8) and data science (49), and $(b)$ the collection of interval decompositions produced by MMA obtained by ranging over all compatible matching functions is itself a topological invariant of the module.
@@
Section 2 provides a concise review of multi-parameter persistence modules and related notions. In Section 3, we present our new descriptors for multi-parameter persistence modules, as well as the MMA algorithm for computing them. In Section 4, we present some approximation properties satisfied by MMA. Then, we provide stronger robustness guarantees for interval decomposable modules in Section 5, and we discuss the design of matching functions in Section 6. Finally, we illustrate the performances of MMA in Section 7.